Assuming that the samples come from normal distribution, find the margin of error E given the following:


n = 10 and x̄ = 28 with s = 4.0, 90% confidence


1
Expert's answer
2021-05-04T13:17:51-0400

n=10, "\bar{x}" =28, s=4, "\alpha" =100-90=10%=0.1


"Z_{\frac{\alpha}{2}}=Z_{0.05}=2.575"


Therefore, The critical value is 2.575.


The standard deviation is 2.0, but as this is a sample, we need the standard error for the mean.

 

The formula for the SE of the mean "\sigma= \dfrac{s}{\sqrt{n}}" ,

so: "\sigma=\dfrac{2}{\sqrt{10}}=0.6325."


Margin of Error="2.575 \times 0.6325 = 1.629."

Therefore, the margin of error is 1.629.